OpenStudy (anonymous):

solve 4(2y-5)is greater than or equal to 3(5y+8)

OpenStudy (matt101):

4(2y-5) >= 3(5y+8) 8y-20 >= 15y+24 -44 >= 7y -44/7 >= y